# On the Identification of Latent Objectives in Stochastic Control

### v1

@@ III Connection with stochastic optimal transport
Consider $\mathcal{C}=\left\{g\in C_{b}(\mathbb{R}^{d}):\|g\|_{BL}\leq 1\right\}$, where $\|g\|_{BL}=\sup_{x\in\mathbb{R}^{d}}|g(x)|+\sup_{x\neq y}|g(x)-g(y)|/|x-y|$. By the well-known density arguments, we have
$\mathcal{D}=(|u|^{2},0)+\bigcup_{\lambda>0}\lambda\tilde{\mathcal{D}}$
with $\tilde{\mathcal{D}}=\{(0,g):g\in\mathcal{C}\}$. Thus (B1) is satisfied. Moreover,
$\rho(\pi,\nu)=\sup_{g\in\mathcal{C}}\,\langle g,\tilde{\pi}_{T}-\tilde{\nu}_{T}\rangle,$
whence $\rho$ defines a metric on $\mathcal{P}(\mathbb{R}^{d})$. More precisely, $\rho(\pi,\nu)=0$ if and only if $\tilde{\pi}_{T}=\tilde{\nu}_{T}$ (see, e.g., Dudley [6]). Consequently, our optimal transport problem becomes
(III.3) $\inf_{u\in\mathcal{U}_{\rho,0}}J(u)=\inf\left\{\mathbb{E}\int_{0}^{T}|u_{t}|^{2}dt:X_{T}^{u}\sim\tilde{\mu}_{T},\;u\in\mathcal{U}\right\},$
@@
Now, using Theorem III.2 we shall derive the duality formula for the Schrödinger’s problem, which is obtained in Mikami and Thieullen [15]. See Mikami [13], [14] for further generalizations. It should be emphasized that our approach is completely different from ones used in [15], [13], and [14]. Recall from Section II that we have assumed
@@
$\sup_{g\in C_{b}(\mathbb{R}^{d})}\left\{\inf_{u\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{1}|u_{t}|^{2}dt+g(X_{1}^{u})\right]-\langle g,\tilde{\mu}_{1}\rangle\right\}$
$=\inf\left\{\mathbb{E}\int_{0}^{1}|u_{t}|^{2}dt:X_{1}^{u}\sim\tilde{\mu}_{1},\;u\in\mathcal{U}\right\}.$
Sketch of the proof.
Applying classical results in theory of the Schördinger problem, we can show that there exists $\{u_{t}^{*}\}_{0\leq t\leq T}\in\mathcal{U}$ such that $\inf_{u\in\mathcal{U}_{\rho,0}}J(u)=J(u^{*})$, whence $u^{*}$ is optimal to the Schrödinger’s problem (III.3). See, e.g. Nutz [20] and Jamison [9].
@@
(III.4) $\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u)=\inf_{u\in\mathcal{U}_{\rho,0}}J(u).$
Let $\varepsilon_{n}$ be a positive sequence such that $\varepsilon_{n}\searrow 0$, and $u^{(n)}\in\mathcal{U}_{\rho,\varepsilon_{n}}$ such that $J(u^{(n)})\leq\inf_{u\in\mathcal{U}_{\rho,\varepsilon_{n}}}J(u)+\varepsilon_{n}$. Put $X^{(n)}=X^{u^{(n)}}$. By a routine argument in probability theory we can show that the sequence $\mathbb{P}(X^{(n)})^{-1}$, $n\in\mathbb{N}$, is tight. So there exist a subsequence $\{n_{k}\}$ and $\hat{P}$ such that $\mathbb{P}(X^{(n_{k})})^{-1}$ weakly converges to $P^{*}$, a path measure of a unique solution of the Schördinger problem. This means that each subsequence $\{\mathbb{P}(X^{n_{k}})^{-1}\}$ contain a further subsequence $\{\mathbb{P}(X^{n_{k_{j}}})^{-1}\}$ converging weakly to $P^{*}$. Applying Theorem 2.6 in [4], we deduce that $\{\mathbb{P}(X^{n_{k}})^{-1}\}$ converges weakly to $P^{*}$ as $n\to\infty$. By the lower semi-continuity of the relative entropy,
$J(u^{*})=\lim_{n\to\infty}J(u^{(n)})\leq\lim_{n\to\infty}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u).$
So (III.4) holds. This together with our theorem III.2 concludes. ∎
@@ Acknowledgements
This study is supported by JSPS KAKENHI Grant Number JP17K05359.

### v2

@@ 3 Connection with stochastic optimal transport
Set $f_{0}(t,x,u)=|u|^{2}$, $g_{0}=0$, and let
$\tilde{\mathcal{D}}=\{(0,g):g\in\mathcal{C}\},$
where $\mathcal{C}=\left\{g\in C_{b}(\mathbb{R}^{d}):\|g\|_{BL}\leq 1\right\}$ with $\|g\|_{BL}=\sup_{x\in\mathbb{R}^{d}}|g(x)|+\sup_{x\neq y}|g(x)-g(y)|/|x-y|$. We apply Theorem 3.2 with $\mathcal{D}$ replaced by $(f_{0},g_{0})+\bigcup_{\lambda>0}\lambda\tilde{\mathcal{D}}\subset\mathcal{D}$. Since $f_{0}=|u|^{2}\geq 0$, the condition (B1) is satisfied: $f_{0}$ is bounded below (hence the proof of Theorem 3.2 applies), $g_{0}=0\in C_{b}(\mathbb{R}^{d})$, and the class $\{\phi_{f,g}:(f,g)\in\tilde{\mathcal{D}}\}$ is non-empty, convex, symmetric at origin, and compact in the topology of uniform convergence on compact subsets of $\mathbb{R}^{d}$. By the Kantorovich–Rubinstein theorem (see, e.g., Dudley [6]),
$\rho(\pi,\nu)=\sup_{g\in\mathcal{C}}\,\langle g,\tilde{\pi}_{T}-\tilde{\nu}_{T}\rangle$
is the bounded-Lipschitz metric on $\mathcal{P}(\mathbb{R}^{d})$. In particular, $\rho(\pi,\nu)=0$ if and only if $\tilde{\pi}_{T}=\tilde{\nu}_{T}$. Consequently, our optimal transport problem becomes
(3.3) $\inf_{u\in\mathcal{U}_{\rho,0}}J(u)=\inf\left\{\mathbb{E}\int_{0}^{T}|u_{t}|^{2}dt:X_{T}^{u}\sim\tilde{\mu}_{T},\;u\in\mathcal{U}\right\},$
@@
Now, using Theorem 3.2 we shall derive the duality formula for the Schrödinger’s problem, which is obtained in Mikami and Thieullen [15]. See Mikami [13], [14] for further generalizations. It should be emphasized that our approach is completely different from ones used in [15], [13], and [14]. Recall from Section 2 that we have assumed
@@
$\sup_{g\in C_{b}(\mathbb{R}^{d})}\left\{\inf_{u\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{T}|u_{t}|^{2}dt+g(X_{T}^{u})\right]-\langle g,\tilde{\mu}_{T}\rangle\right\}=\inf\left\{\mathbb{E}\int_{0}^{T}|u_{t}|^{2}dt:X_{T}^{u}\sim\tilde{\mu}_{T},\;u\in\mathcal{U}\right\}.$
Proof.
Let $p(t,x,s,y)$ be the transition density of $d$-dimensional Brownian motion. Then, there exists a $\sigma$-finite product measure $\nu_{0}(dx)\nu_{T}(dy)$ such that
$\pi(E):=\int_{E}p(0,x,1,y)\nu_{0}(dx)\nu_{T}(dy),\quad E\in\mathcal{B}(\mathbb{R}^{d}\times\mathbb{R}^{d}),$
satisfies $\pi(dx\times\mathbb{R}^{d})=\tilde{\mu}_{0}(dx)$ and $\pi(\mathbb{R}^{d}\times dy)=\tilde{\mu}_{T}(dy)$ and
$H(\pi\;|\;p(0,x,1,y)\tilde{\mu}_{0}(dx)dy)<\infty,$
where $H(\pi\,|\,R)$ denotes the relative entropy of $\pi$ with respect to $R$. See e.g., Theorem 2.1 in Nutz [20]. Then, by Theorem 2 in Jamison [9], there exists a weak solution of
$dX_{t}^{*}=\nabla h(t,X_{t}^{*})dt+dB_{t},\quad X_{0}^{*}\sim\tilde{\mu}_{0},$
where
$h(t,x)=\int_{\mathbb{R}^{d}}p(t,x,1,y)\nu_{T}(dy),\quad(t,x)\in[0,T)\times\mathbb{R}^{d}.$
We assume that our probability space, filtration, and Brownian motion are identical to those used to construct this weak solution. Then, the control process $\{u_{t}^{*}\}_{0\leq t\leq T}$ defined by $u^{*}_{t}=\nabla h(t,X_{t}^{*})$ is in $\mathcal{U}$ and satisfies
$\inf_{u\in\mathcal{U}_{\rho,0}}J(u)=J(u^{*})=2H(\pi\;|\;p(0,x,1,y)\tilde{\mu}_{0}(dx)dy),$
whence $u^{*}$ is optimal to the Schrödinger’s problem (3.3). In particular, the finiteness condition (3.1) is satisfied.
@@
(3.4) $\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u)=\inf_{u\in\mathcal{U}_{\rho,0}}J(u).$
To this end, let $P^{W},P^{*}\in\mathcal{P}(\mathbb{W}^{d})$ be the law of $X_{0}+B$ and $X^{*}$, respectively. Then it is well-known that
(3.5) $\inf H(Q\,|\,P^{W})=H(P^{*}\,|\,P^{W}),$
where the infimum is taken over all $Q\in\mathcal{P}(\mathbb{W}^{d})$ with $Q_{0}=\tilde{\mu}_{0}$ and $Q_{T}=\tilde{\mu}_{T}$. Here $Q_{0}$ and $Q_{T}$ denote the marginals of $Q$ at time 0 and $T$, respectively. Let $\varepsilon_{n}$ be a positive sequence such that $\varepsilon_{n}\searrow 0$. Let $u^{(n)}\in\mathcal{U}_{\rho,\varepsilon_{n}}$ be such that
$J(u^{(n)})\leq\inf_{u\in\mathcal{U}_{\rho,\varepsilon_{n}}}J(u)+\varepsilon_{n}.$
Put $X^{(n)}=X^{u^{(n)}}$. Let $\delta>0$ be arbitrary. For a fixed $N\geq 1$ take $\{t_{i}\}$ so that $0=t_{0}<t_{1}<\cdots<t_{N}=T$ and $t_{\ell}-t_{\ell-1}=T/N$. Then from Theorem 7.4 in Billingsley [4] it follows that
$\mathbb{P}\left(\max_{\begin{subarray}{c}0\leq s<t\leq T\\ 0<t-s<T/N\end{subarray}}|X_{t}^{(n)}-X_{s}^{(n)}|>\delta\right)$
$\leq\mathbb{P}\left(\max_{\begin{subarray}{c}0\leq s<t\leq T\\ 0<t-s<T/N\end{subarray}}\left|\int_{s}^{t}u_{r}^{(n)}dr\right|>\delta/2\right)+\sum_{\ell=1}^{N}\mathbb{P}\left(\sup_{t_{\ell-1}\leq s\leq t_{\ell}}|B_{s}-B_{t_{\ell-1}}|\geq\delta/6\right).$
The first term of the right-hand side is at most
$\frac{4}{\delta^{2}N}\sup_{n}\mathbb{E}\int_{0}^{T}\left[1+|u_{r}^{(n)}|^{2}\right]dr\leq\frac{4}{\delta^{2}N}\left[T+\sup_{n}\mathbb{E}\int_{0}^{T}|u_{r}^{(n)}|^{2}dr\right].$
The second term is at most $C_{0}\delta^{-3}N^{-1/2}$ where we have used Doob’s maximal inequality. Thus, for any $\delta>0$,
$\lim_{N\to\infty}\sup_{n\geq 1}\mathbb{P}\left(\max_{\begin{subarray}{c}0\leq s<t\leq T\\ 0<t-s<T/N\end{subarray}}|X_{t}^{(n)}-X_{s}^{(n)}|>\delta\right)=0.$
Therefore, by Theorem 7.3 in Billingsley [4], the sequence of laws of $X^{(n)}$, $n\in\mathbb{N}$, is tight. Denote by $Q^{(n)}$ the probability distribution of $X^{(n)}$ under $\mathbb{P}$. So there exist a subsequence $\{n_{k}\}$ and $\hat{P}\in\mathcal{P}(\mathbb{W}^{d})$ such that $Q^{(n_{k})}$ weakly converges to $\hat{P}$. Since $\rho$ is a metric on $\mathcal{P}(\mathbb{R}^{d})$ we have
$\rho(\hat{P},\tilde{\mu})\leq\rho(\hat{P},Q^{(n_{k})})+\rho(Q^{(n_{k})},\tilde{\mu})\to 0,$
as $k\to\infty$, whence $\hat{P}_{1}=\tilde{\mu}_{1}$. Further, by the lower semi-continuity of $Q\mapsto H(Q|P)$,
$H(P^{*}\,|\,P^{W})\leq H(\hat{P}\,|\,P^{W})=\liminf_{k\to\infty}H(Q^{(n_{k})}\,|\,P^{W})=\frac{1}{2}\liminf_{k\to\infty}J(u^{(n_{k})})\leq H(P^{*}\,|\,P^{W}).$
This means that $\hat{P}$ is optimal to (3.5), whence by uniqueness, we obtain $\hat{P}=P^{*}$. What we have shown now is that each subsequence $\{Q^{(n_{k})}\}$ contains a further subsequence $\{Q^{(n_{k_{j}})}\}$ converging weakly to $P^{*}$. Applying Theorem 2.6 in [4], we deduce that $\{Q^{(n_{k})}\}$ converges weakly to $P^{*}$ as $k\to\infty$. Again by the lower semi-continuity of the relative entropy,
$\liminf_{n\to\infty}J(u^{(n)})=2\liminf_{n\to\infty}H(Q^{(n)}\,|\,P^{W})\geq 2H(P^{*}\,|\,P^{W})\geq\limsup_{n\to\infty}J(u^{(n)}).$
Hence,
$J(u^{*})=2H(P^{*}\,|\,P^{W})=\lim_{n\to\infty}J(u^{(n)})\leq\lim_{n\to\infty}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u).$
So (3.4) holds. This together with Theorem 3.2 concludes.
Conclusion
We have studied the inverse stochastic control problem from a variational perspective. Given an observed state-control distribution $\{\mu_{t}\}$, which need not be realizable as the marginal law of any controlled process, we introduced the suboptimality gap functional $V(f,g)$ as a natural criterion for identifying latent cost functions. Under mild regularity conditions, we established the existence of a minimizer for the resulting inverse problem (1.4) (Proposition 2.3).
The main contribution of the paper is a variational equivalence between the inverse problem and a generalized dynamic Schrödinger problem (Theorem 3.2): the optimal value $V^{*}$ coincides with the value of a stochastic control problem with relaxed marginal constraints, measured by the pseudometric $\rho$ induced by the class $\tilde{\mathcal{D}}$ of latent objectives. As a consequence, the classical duality formula for the Schrödinger bridge problem is recovered as a special case (Corollary 3.4), via a different approach from those in [15, 13, 14].
Several directions remain open. A regularization of the inverse problem, which would replace the hard constraint $\rho(P^{u},\mu)=0$ by a penalized formulation, is a natural extension. Uniqueness of the minimizer under general conditions, and computational methods for solving the inverse problem in practice, are also subjects of future work.
@@ Acknowledgements
This study is supported by JSPS KAKENHI Grant Number JP24K06861.